$SPAD/src/input e01baf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
m:=7
x:Matrix SF:=
 [[0.0 ,0.2 ,0.4 ,0.6 ,0.75 ,0.9 ,1.0 ]]
y:Matrix SF:=
 [[1.0000 ,1.2214 ,1.4918 ,1.8221 ,2.1170 ,2.4596 ,2.7183 ]]
lck:=11
lwrk:=58
result:=e01baf(m,x,y,lck,lwrk,-1)

———-

References^{[1] nothing}

4